a) A metal ring of 2.40 cm diameter moves into the magnetic field of an MRI scanner in 0.30 s from where the magnetic field is zero. The magnetic field in the MRI is 1.90 T and the plane of the ring is perpendicular to this field. Determine the voltage induced in the ring.

b) If the ring in the problem above has a resistance of 0.009 Ω, what is the current induced in the ring, and in what direction (looking into the MRI)? What is the value of the magnetic field induced at the center of the ring?
Answer #1

R= Radive of ring =02(m=1,2x10m liato 419 T It = 1 = 1 RAB FR 다 À & At volt 3 47 a> AB 6-3 3 S indured Embze = AAB At =(1-2)xMagnetic held indured at the center of the ring, Вс - Ho T 45 р 1 -) x0-32x2л 2 ) 2 —S - - 67 x) T - S А. — Вс = 167ѕ хірт зо